If x=5y=2z then find x:y:z

Answer by Theo(13342)

what i get is the following:

start with x=5y=2z

if we put everything in terms of z, then we get:

x = 2z

y = 2/5 * z

x:y:z in terms of z is therefore equalto 2z:2/5z:z

multiply everything by 5 and you get 10z:2z:5z

divide everything by z and you get 10 : 2 : 5

the ratio of x : y : z is equal to 10 : 2 : 5

there are 10 x's for every 2 y's for every 5 z's.

the ratio of x/y = 10/2 = 5/1

solve for x and you get x = 5y.

the ratio of y/z = 2/5

cross multiply to get 5y = 2z.

the ratio of x : y : z looks good at 10 : 2 : 5